OpenStudy (anonymous):

What's answer of (sin 160 + sin 20) / (cos 140+cos 70)

OpenStudy (debbieg):

Are you supposed to use a calculator? Those aren't common unit circle angle measures and I don't see a way to simplify with any identities.... so I think you would just punch it up in your calculator.

OpenStudy (anonymous):

I have calculated that (sin 160 + sin 20) is 2 sin 20. But, I am so confused about ((cos 140+cos 70)....

OpenStudy (debbieg):

Well, using cos(x)=sin(90-x) we can get that cos(70)=sin(20) And cos(140)=-cos(40)=-cos(2*20) So (cos 140+cos 70)=-cos(2*20)+sin(20) so that makes the whole thing: (sin 160 + sin 20) / (cos 140+cos 70)=2*sin(20)/(sin(20)-cos(2*20)) So that puts everything in terms of 20*. But I'm still not sure where to go with it after that, if you need an exact numerical evaluation...

OpenStudy (anonymous):

I have calculated what your answer above, and I am also confused... DebbieG ...

OpenStudy (mertsj):

\[\sin 160 + \sin 20 = 2\sin 90\cos 70\] \[\cos 140 + \cos 70 = 2\cos 105\cos 35\]

OpenStudy (anonymous):

Yes, Mrtsj, thanks beforehand... I have reached what you write, but I am confused to continue it...

OpenStudy (mertsj):

Do the directions say to simplify or what?

OpenStudy (anonymous):

not simplify, but to calculate the final answer...

OpenStudy (anonymous):

I am still confused, there isn't the answer what I mean on the website...

OpenStudy (debbieg):

Ahhh, I had forgotten about sum-to-product identities. But yeah, I'm still not sure how you could evaluate an exact result here without a calculator. I mean, you can certainly simplify: \( \dfrac{\sin 160 + \sin 20}{cos 140 + \cos 70 } =\dfrac{ 2\sin 90\cos 70}{2\cos 105\cos 35}=\dfrac{ \cos 70}{\cos 105\cos 35}\) Hmmmmmm...... now, the only thing I notice about this is that it is of the form: \( \dfrac{ \cos \alpha}{\cos (\alpha + 45)\cos (\alpha/2)}\) I have no idea if that's relevant or how it might help, though... lol.

OpenStudy (anonymous):

.9933763474? idek..

OpenStudy (anonymous):

But i want to calculate it without a calculator....